OpenStudy (chillout):

How many pairs of integers (x,y) are there such as x²y-xy²=3015

OpenStudy (chillout):

That's how I started the question: \(xy(x-y)=3015\\xy\neq0,\\x-y\neq0,\\x\neq0,\\y\neq0\) Given that are there algebraic solutions?

OpenStudy (chillout):

Further factorisation gives \(xy(x-y)=3^{2}*5*67\) I guess it's by trial and error, then?

imqwerty (imqwerty):

i was doing the same thing

imqwerty (imqwerty):

but i guess that would be lengthy :/

imqwerty (imqwerty):

normally i isolate one of the given unknowns and write it in terms of the other unknown and then i use the information that "both x and y are integers" to find the possible combinations

OpenStudy (chillout):

This would be impractical, I guess. I'd give a try doing something about the prime factorization and the given conditions. But dunno where to start :(

OpenStudy (chillout):

Well, I have the alternatives though: (A) 1005 (B) 603 (C) 8 (D) 1 (E) 0

imqwerty (imqwerty):

yeah isolating them seems impractical i'm trying to do some useful substitutions for x and y like rn i'm trying to convert it into a useful form by substituting x=t+1 and y=p+1

OpenStudy (chillout):

http://www.wolframalpha.com/input/?i=solve+x%C2%B2y-xy%C2%B2%3D3015+over+integer+values. It uses quadratic formula/completing the square. But the answer is not convicing

imqwerty (imqwerty):

yeah i totally forgot that thing uh we can assume it to be a quadratic in y and then by using quadratic formula you can get the values of y remember that y is an integer so the discriminant must be a perfect square so using this you get the solution pairs

ganeshie8 (ganeshie8):

Hint : x and y must be odd

OpenStudy (chillout):

Yeah, given the factors.

ganeshie8 (ganeshie8):

xy(x-y) = 3015 Since 3015 is odd, we can say x, y, and x-y must be odd

ganeshie8 (ganeshie8):

When x and y are odd, what can we say about x-y ?

OpenStudy (chillout):

There's no solutions for x-y being odd. Now that's a simple way!

OpenStudy (chillout):

This proof is good enough. xy will be odd, but x-y will be always even. GIven that, xy*(x-y) will always be even.

ganeshie8 (ganeshie8):

Looks nice and simple to me!

OpenStudy (chillout):

Thanks, saved me a bunch of impractical stuff :D

OpenStudy (agent0smith):

@ChillOut from this you should learn to always trust wolframalpha ;) Good question. Nice job @ganeshie8

OpenStudy (chillout):

@agent0smith I trust it and I know it's right. Hell, I use it a lot! I meant It wasn't a proof that should be used :D

OpenStudy (agent0smith):

haha i know what you meant :D